OpenStudy (anonymous):

evaluate tan(arccos(2/5))

OpenStudy (anonymous):

 there is a picture of an angle whose cosine is \(\frac{2}{5}\) what you need is the "opposite side" which you find via pythagoras

OpenStudy (dumbcow):

\[\tan = \sqrt{\sec^2 x - 1} = \sqrt{\frac{1}{\cos^2 x} -1}\]

OpenStudy (anonymous):

solve \[2^2+b^2=5^2\] for \(b\) and then your answer will be \(\frac{b}{5}\)